Spy Agent 001 is chasing the bad guys. Suddenly, 001 needs to turn his car (m = 1000 kg) around, and completes a turn in the shape of a semi-circle (radius = 60 m) travelling at a constant speed of 27 m/s.

a) If the static coefficient of friction between the tires and the pavement is 0.55, find the maximum force of friction before slipping occurs.

b) Will the car be able to complete the turn without sliding?

c) To complete the turn safely (without sliding), what is the minimum turning radius required if the speed of the car is 27 m/s?

d) To complete the turn safely (without sliding), what is the maximum speed if the turning radius is 60 m?
Answer #1

Slipping occurs at

v = sqrt(us g R)

v = 18.0 m/s   [ANSWER, PART D, I did this first]

At maximum frictional force is

F = us m g = 5390 N   [ANSWER, PART A]

****************

As 27 m/s is greater than the critical speed at part A, no, he won't make it. [ANSWER, PART B]

***************

v = sqrt(us g R)

--> R = v^2 / [us g]

R = 135 m   [ANSWER, PART C]

**************

Slipping occurs at

v = sqrt(us g R)

v = 18.0 m/s   [ANSWER, PART D]